OpenStudy (anonymous):

2^(x+1) = 3^(x-1) could someone explain in detail how to solve these type of questions please?

OpenStudy (anonymous):

Become friendly with logarithms. Because \[\log(a^b) = b\cdot \log(a)\] they are very useful in situations like this. For instance, taking the log of both sides, \[\log(2^{x+1}) = (x+1)\log(2) = \log(3^{x-1}) = (x-1)\log(3)\] Pulling the x's to one side and the constants to the other, we find \[\left(\log(2)-\log(3)\right)x = -\left(\log(3)+\log(2)\right) \] so we get \[ x= \frac{\log(3)+\log(2)}{\log(3)-\log(2)} \approx 4.42 \]

OpenStudy (anonymous):

As a side note, it doesn't actually matter which type of logarithm you use. Base 10, base e, base 2, it all works equally well.

OpenStudy (anonymous):

how did you get log(2)-log(3))x= - (log(3)+log(2)) ?

OpenStudy (anonymous):

Just rearrange the equation. Get all the x-terms on one side, and the constants on the other.

OpenStudy (anonymous):

so first its like (x+1)log2= (x-1)log 3, right?

OpenStudy (anonymous):

right... then expand that, to get \[x\log(2) + \log(2) = x\log(3) - \log(3)\]

OpenStudy (anonymous):

thanks !